OpenStudy (pannda_lover4):

After Elise deposits $21.75 into her bank account, her balance is $285.00. Which equation can be used to find the amount n in her account before she made the deposit? Select one: a. n + 285.00 = 21.75 b. n – 285.00 = 21.75 c. 21.75 + n = 285.00 d. 21.75 – n = 285.00

OpenStudy (pannda_lover4):

will fan and medal

OpenStudy (legomyego180):

What do you have to do to figure out how much she had before?

OpenStudy (legomyego180):

In other words, if I said hey I just deposited 21.75 to my bank account, now I have $285 dollars but I need to know how much I had before. How would you figure that out?

OpenStudy (pannda_lover4):

by subtracting @legomyego180

OpenStudy (legomyego180):

yea so you could say: n = the amount you had before \[n=285-21.75\] Notice thats not one of your answer choices so you can rewrite this a bit by adding the 21.75 to the other side of the equation like this: \[n+21.75=285\]

OpenStudy (pannda_lover4):

@legomyego180 That is not on of the choices

OpenStudy (legomyego180):

If you care about it: they try to trick you with the first two answers. You could also rewrite this as: \[n-285=-21.75\]

OpenStudy (legomyego180):

yes it is, its C 21.75 + n = n +21.75

OpenStudy (pannda_lover4):

ok thanks

OpenStudy (bondjames):

oh i was too late